8. Suppose a ball is drawn at random from a box containing three white and three black balls. After a ball is drawn, it is then replaced and another drawn. What is the probability that of the first four balls drawn, exactly two are white?

9. In a bolt manufacturing factory, machines A, B, and C produce 25%, 30% and 45% of the total output, respectively. Of their outputs, 7%, 6% and 4% are defective bolts,
respectively. If a bolt drawn at random from the production is found to be defective,
what is the probability that it was manufactured by machine C?
1
Expert's answer
2020-06-01T18:21:00-0400

8.


"P(WWBB)+P(WBBW)+P(WBWB)+""+P(BWWB)+P(BWBW)+P(BBWW)="

"=({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})+({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})+"


"+({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})+({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})+"

"+({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})+({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})+"

"+({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})+({3 \over 6})({3\over 6})({3 \over 6})({3 \over 6})="

"=6({1\over 16})={3\over 8}"

9. Let 

"E_1:" the event that bolt is produced by machine A,

"E_2:" the event that bolt is produced by machine B, and 

"E_3:" the event that bolt is produced by machine C.

Here "E_1, E_2," and "E_3" are mutually exclusive and exhaustive events.

We have


"P(E_1)=0.25,P(E_2)=0.3,P(E_3)=0.45"

Let "D:" the event that bolt chosen is found to be defective.

Then


"P(D|E_1)=0.07,P(D|E_2)=0.06,P(D|E_3)=0.04"

From Bayes' Theorem


"P(E_3|D)={P(D|E_3)P(E_3)\over P(D|
E_1)P(E_1)+P(D|E_2)P(E_2)+P(D|E_3)P(E_3) }="

"={0.04(0.45)\over 0.07(0.25)+0.06(0.3)+0.04(0.45) }={36\over 107}\approx0.3364"